OpenStudy (anonymous):

Ok, here's a crazy one. Melinda and JP are selling Lemonade at a stand in her front yard. The linear expression below models her total profit in dollars after h hours. 2 - (6h-5)+4h 3 a. What is her profit after 2 hours? Explain b. When will her profit be $30? Explain c. How much does her profit increase each hour? Explain d. Evaluate the expression for h=0. What does this value mean in the context of the problem? e. Explain how the answer to part 3 could be calculated mentally just by looking at the expression. f. Explain how the answer to part 4 could be calculated mentally just

OpenStudy (cruffo):

Looks like you could simplify your formula a little \[\frac{2}{3}(6h- 5)+4h \] \[= 4h - \frac{10}{3} + 4h\] \[ = 8h - \frac{10}{3} \] We could go with decimals, but since the question uses fractions, I would continue to use fractions.

OpenStudy (cruffo):

(a) They are giving you 2 hours and asking for profit - this means that they have given you an h value and they want you to find a ... p ??? value. Substitute 2 for h. \[p = 8(2) - \frac{10}{3} = \; ...\] (b) They are telling you $30 profit and want you to find hours - this means that they have given you a p value and want you to solve for h. Set the formula equal to 30 and solve for h. \[30 = 8h - \frac{10}{3} \]

OpenStudy (anonymous):

I'm with you so far :o)

OpenStudy (cruffo):

ok! (c) the equation is that of a line, so the are asking you about the slope of the line. (d) When h = 0, this is at the beginning, before any time has passed. Graphically, it's the y-intercept of the line.

OpenStudy (cruffo):

(e) y= mx + b, just look for the number with a variable, that's the slope of the line. (f) y = mx + b, just look for the number without a variable, that's the y-intercept

OpenStudy (anonymous):

Absolutely fabulous! Thank you for your time and efforts in this problem. I really appreciate it!!